The steam turbine for a new biomass power plant in Sarawak operates at 600C at 7000kPa. The exhaust steam exit the turbine at 50C with 4% (by mass) of the steam is condensed to liquid. Neglecting pump work and kinetic and potentialenergy changes, calculate the thermal efficiency of the cycle and the turbine efficiency. Please state your assumptions accordingly

.Boiler turbine Turbine pump) Condenser he
Answer #1

Solution → 은 dan P=7000 KP , T = 600°C Boiler h = 3650-3 K5/89 2 1 $2 = 7:0894 KJ / kgK Wis (turbine) Turbine 7. this from P=Shower # ne Toshine = W W Ws hy-hy) thy-hy) +(3650-3–2496-702) Katig (3650-3-2273-20 208) Koky # Me Twiline = 0.8377 8% Musli

In case of any doubt or you need any further help regarding this solution please comment below.

Thank you. Please upvote it.

Please...........